OpenStudy (anonymous):

Solve for x: 3x2 + 7x = 2

OpenStudy (anonymous):

6+7x=2 7x=-4 x=-4/7

OpenStudy (anonymous):

set it =0 \[3x ^{2}-7x+2=0\] a=3 b=-7 c=2 the quadratic formula=\\[=-b \pm \sqrt{b^2-4(a)(c)}\div2(a)\]